Determine the mean, median, and mode of following data set: 83 78.8 98 57.1 98 77.4 67.8 98 61.9 70.2 Mean- Median- Mode- Round your answers to the nearest hundredth as needed Get help: Video Points possible: 3 This is attempt 1 of 1. Submit
Answer #1

Solution:

Given data set
83, 78.8, 98, 57.1, 98, 77.4, 67.8, 98, 61.9, 70.2

Sum of terms = 83 + 78.8 + 98 + 57.1 + 98 + 77.4 + 67.8 + 98 + 61.9 + 70.2 = 790.2
Number of terms = 10
Mean = sum of terms/Number of terms = 790.2/10 = 79.02

The median is the middle number in a sorted list of numbers. So, to find the median, we need to place the numbers in value order and find the middle number Ordering the data from least to greatest, we get: 57.1 61.9 67.8 70.2 77.4 78.8 83 98 98 98 As you can see, we do not have just one middle number but we have a pair of middle numbers, so the median is the average of these two numbers 77.4+78.8 -78.1 Median - 78.1

The mode of a set of data is the value in the set that occurs most often.

Ordering the data from least to greatest, we get:

57.1 61.9 67.8 70.2 77.4 78.8 83 98 98 98   

We see that the mode is 98.